# Wasserstein distances and divergences of order $p$ by quantum channels

### v2

@@ 3. $p$-Wasserstein distances and divergences
where $\Pi_{*}^{(\alpha,\beta)}$ stands for a $C_{\mathcal{A},p}$-optimal coupling of the states $\alpha$ and $\beta.$ Although the definition (25) is highly natural in view of the definition of the quadratic Wasserstein divergence [17], numerical tests indicate that
@@
may happen for $p>2$ and hence the right-hand side of (25) is ill-defined in this case. Indeed, if $\mathcal{H}=\mathbb{C}^{3},$ and the states $\rho_{0},\omega_{0}\in\mathcal{S}(\mathbb{C}^{3})$ are given in the computational basis by
$\rho_{0}=\left[\begin{array}[]{ccc}0.317&-0.219-0.299i&0.177-0.028i\\ -0.219+0.299i&0.507&-0.049+0.241i\\ 0.177+0.028i&-0.049-0.241i&0.176\end{array}\right]$ (27)
and
$\omega_{0}=\left[\begin{array}[]{ccc}0.415&0.112+0.081i&0.365+0.105i\\ 0.112-0.081i&0.153&0.164-0.102i\\ 0.365-0.105i&0.164+0.102i&0.432\end{array}\right],$ (28)
and
$\mathcal{A}_{0}=\left\{\left[\begin{array}[]{ccc}-2.991&-0.119+1.802i&1.033-3.505i\\ -0.119-1.802i&-2.806&1.300+3.082i\\ 1.033+3.505i&1.300-3.082i&-0.370\end{array}\right]\right\},$ (29)
that is, the transport cost is generated by a single observable given above, then for $p_{0}=2.35$ we get
$D_{\mathcal{A}_{0},p_{0}}^{p}(\rho_{0},\omega_{0})-\frac{1}{2}\left(D_{\mathcal{A}_{0},p_{0}}^{p}(\rho_{0},\rho_{0})+D_{\mathcal{A}_{0},p_{0}}^{p}(\omega_{0},\omega_{0})\right)=-0.572<0.$ (30)
@@ 4. Relations between the sets of all possible cost operators for different values of $p$
$\mathbf{C}_{\infty}(\mathcal{H})=\{C\geq 0\;|\;\forall\epsilon>0,\exists p^{\prime}:\forall p\geq p^{\prime},\exists C^{\prime}\in\mathbf{C}_{p}(\mathcal{H}):\left|\left|C-C^{\prime}\right|\right|_{op}\leq\epsilon\}$ (32)
@@
So let $0<p<\infty$ and $A\in\mathcal{L}(\mathcal{H})^{sa},\,\#\{\mathrm{spec}(A)\}<\infty$ be fixed, and assume that $A$ admits the spectral decomposition $A=\sum_{r=1}^{n}\lambda_{r}Q_{r}.$ Let $J\subseteq[n]$ where the shorthand $[n]:=\left\{1,\dots,n\right\}$ is used. Let us define
@@
$=\left|\alpha_{J}\right|^{p}\sum_{r\in[n]}\sum_{s\in[n]}\left|\mathbf{1}_{\left\{r\in J\right\}}-\mathbf{1}_{\left\{s\in J\right\}}\right|Q_{r}\otimes Q_{s}^{T}.$ (37)
@@
holds for every $r,s\in[n].$ The function $[n]\times[n]\ni(r,s)\mapsto\left|\lambda_{r}-\lambda_{s}\right|$ appearing on the left-hand side of (41) is clearly an $l_{1}$-embeddable semi-metric, that is, there is a positive integer $m$ and points $u_{1},\dots,u_{n}\in\mathbb{R}^{m}$ such that $\left|\lambda_{r}-\lambda_{s}\right|=\left|\left|u_{r}-u_{s}\right|\right|_{l_{1}(\mathbb{R}^{m})}.$ The trivial choice $m=1$ and $u_{r}:=\lambda_{r}$ for $r\in\left\{1,\dots,n\right\}$ does the job. Therefore, by Proposition 4.2.2. of [18], the function $(r,s)\mapsto\left|\lambda_{r}-\lambda_{s}\right|$ belongs to the cut cone $\mathrm{CUT}_{n}$. By the definition of the cut cone, see Eq. (4.1.2.) of [18], this means that there is an element-wise non-negative vector $\left\{\beta_{J}\right\}_{J\subseteq[n]}$ such that
@@
If $A\in\mathcal{L}(\mathcal{H})^{sa}$ and #$\#\{\mathrm{spec}(A)\}\leq 2,$ then $A=\lambda Q+\lambda^{\prime}(I-Q)$ for some $Q\in\mathcal{P}(\mathcal{H})$ and $\lambda,\lambda^{\prime}\in\mathbb{R}.$ Therefore,
$\left|A\otimes I^{T}-I\otimes A^{T}\right|^{p}=\left|\lambda Q\otimes Q^{T}+\lambda^{\prime}(I-Q)\otimes Q^{T}+\lambda Q\otimes(I-Q)^{T}+\lambda^{\prime}(I-Q)\otimes(I-Q)^{T}-\right.$
@@
$=\left|\lambda-\lambda^{\prime}\right|^{p}\left((I-Q)\otimes Q^{T}+Q\otimes(I-Q)^{T}\right).$
@@
$=\left|\mu\right|^{p^{\prime}}\left((I-Q)\otimes Q^{T}+Q\otimes(I-Q)^{T}\right).$
@@
$\mathbf{C}_{p}(\mathcal{H})\subseteq\mathbf{C}_{1}(\mathcal{H})\subseteq\mathbf{C}_{p^{\prime}}(\mathcal{H})\text{ and }\mathbf{C}_{p^{\prime}}(\mathcal{H})\subseteq\mathbf{C}_{1}(\mathcal{H})\subseteq\mathbf{C}_{p}(\mathcal{H}).$ (54)
@@
Let $\mathcal{H}$ be a separable Hilbert space, then for any $p^{\prime}\geq p>0$ and any $A=\sum_{i=1}^{n}\lambda_{i}P_{i}\neq c\cdot I$ observable on $\mathcal{H}$ there exists an observable $B=\sum_{i=1}^{n}\mu_{i}P_{i}$ and an observable $A^{\prime}=\sum_{i=1}^{n}\lambda^{\prime}_{i}P_{i}$ on $\mathcal{H}$, such that #$\#\left\{\mathrm{spec}\left(A^{\prime}\right)\right\}<\#\left\{\mathrm{spec}\left(A\right)\right\}$ and
@@
The following proposition complements Conjecture 8 and asserts that the map $p\mapsto\mathbf{C}_{p}(\mathcal{H})$ is not only monotone increasing but strictly monotone increasing on $[1,\infty).$
@@
$\left|\lambda_{1}-\lambda_{2}\right|^{p^{\prime}}+\left|\lambda_{2}-\lambda_{3}\right|^{p^{\prime}}-2^{1-\tilde{p}}\left|\lambda_{1}-\lambda_{3}\right|^{p^{\prime}}=1+1-2^{1-\tilde{p}}\cdot 2^{p^{\prime}}=2-2\cdot 2^{(p^{\prime}-\tilde{p})}=2(1-2^{(p^{\prime}-\tilde{p})})$
@@
$\mathbf{C}_{p}(\mathbb{C}^{3})\subseteq\mathbf{C}_{p^{\prime}}(\mathbb{C}^{3})\text{ and }\mathbf{C}_{p}(\mathbb{C}^{3})\nsupseteq\mathbf{C}_{p^{\prime}}(\mathbb{C}^{3}).$
@@
$C_{\{A(p)\},p}=\lambda(Q_{1}\otimes Q_{2}^{T}+Q_{2}\otimes Q_{1}^{T})+$
$+\frac{\lambda}{2^{p}}(Q_{1}\otimes(I-Q_{1}-Q_{2})^{T}+(I-Q_{1}-Q_{2})\otimes Q_{1}^{T}+Q_{2}\otimes(I-Q_{1}-Q_{2})^{T}+(I-Q_{1}-Q_{2})\otimes Q_{2}^{T})$ (78)
and
$\lim_{p\rightarrow\infty}\left|\left|C_{\{A(p)\},p}-\lambda(Q_{1}\otimes Q_{2}^{T}+Q_{2}\otimes Q_{1}^{T})\right|\right|_{op}=0.$ (79)
@@ 5. Triangle inequality is inherited by larger parameter values
$[0,\infty)\ni t\mapsto t^{\frac{\min\left\{1/p^{\prime},1\right\}}{\min\left\{1/p,1\right\}}}$ (88)
is subadditive. Subadditivity follows from concavity and from the fact that it maps 00 to itself. By the triangle inequality (81), the obvious monotonicity of (88) and the subadditivity of (88) we get
@@
$+\left(D_{\mathcal{A},p}^{\max\left\{p,1\right\}}(\tau,\omega)-\frac{1}{2}\left(D_{\mathcal{A},p}^{\max\left\{p,1\right\}}(\tau,\tau)+D_{\mathcal{A},p}^{\max\left\{p,1\right\}}(\omega,\omega)\right)\right)^{\min\left\{\frac{1}{p^{\prime}},1\right\}}$
@@
$+\left(D_{\mathcal{A}^{\prime},p^{\prime}}^{\max\left\{p^{\prime},1\right\}}(\tau,\omega)-\frac{1}{2}\left(D_{\mathcal{A}^{\prime},p^{\prime}}^{\max\left\{p^{\prime},1\right\}}(\tau,\tau)+D_{\mathcal{A}^{\prime},p^{\prime}}^{\max\left\{p^{\prime},1\right\}}(\omega,\omega)\right)\right)^{\min\left\{\frac{1}{p^{\prime}},1\right\}}$
$=d_{\mathcal{A}^{\prime},p^{\prime}}(\rho,\tau)+d_{\mathcal{A}^{\prime},p^{\prime}}(\tau,\omega).$ (90)
@@ 6. $p=2$ as a possible threshold for the triangle inequality
$=\inf_{\Pi\in\mathcal{C}(\rho,\omega)}\left\{\sum_{k=1}^{K}\left(\mathrm{tr}_{\mathcal{H}}\left[\rho A_{k}^{2}+\omega A_{k}^{2}\right]-2\mathrm{tr}_{\mathcal{H}\otimes\mathcal{H}^{*}}\left[\Pi\cdot A_{k}\otimes A_{k}^{T}\right]\right)\right\}.$ (91)
@@
An important feature of the quadratic Wasserstein distances is that the distance of a state $\rho$ from itself (which may be positive) is always realized by the identity channel — see [16, Corollary 1]. Therefore, the quadratic Wasserstein distance of a state $\rho\in\mathcal{S}\left(\mathcal{H}\right)$ from itself admits the explicit form
@@
For every $20<p<$ 2 there is $\mathcal{H}$ and $\mathcal{A}\in\mathbf{FCO}(\mathcal{H})$ and $\rho,\tau,\omega\in\mathcal{S}\left(\mathcal{H}\right)$ such that
@@
and the positivity condition $\rho\geq 0$ is equivalent to the Euclidean length of $\mathbf{b}_{\rho}$ being at most $1$ Now observe that Proposition 16 tells us that
@@ 7. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 19
This section is dedicated to the study of the triangle inequality for quadratic quantum Wasserstein divergences. The main result here is a proof of the triangle inequality using the sole assumption that an arbitrary one of the three states involved is pure. This is a generalization of the main result of [5], where we proved the triangle inequality in the case when the intermediate state is pure.
@@
$+\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\tau^{1/2}A_{k}\right)^{2}+\left(\omega^{1/2}A_{k}\right)^{2}-2\tau^{1/2}A_{k}\tau^{1/2}\Psi_{opt}^{\dagger}\left(A_{k}\right)\right]\right)^{1/2}\geq$
@@
$+\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\tau^{1/2}A_{k}\right)^{2}+\left(\tau^{1/2}\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)^{2}-2\tau^{1/2}A_{k}\tau^{1/2}\Psi_{opt}^{\dagger}\left(A_{k}\right)\right]\right)^{1/2}=$
@@
is a special case of the monotonicity version of Lieb’s concavity — see, e.g., [9, Theorem 1]. We can use (117) for a further estimation in the following way:
@@
$+\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\rho^{1/2}\left(\Phi_{opt}^{\dagger}\left(A_{k}\right)-\Phi_{opt}^{\dagger}\left(\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)\right)\right)^{2}\right]\right)^{1/2}\geq$
@@
$=\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\rho^{1/2}A_{k}\right)^{2}+\left(\rho^{1/2}\Phi_{opt}^{\dagger}\left(\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)\right)^{2}-2\rho^{1/2}A_{k}\rho^{1/2}\Phi_{opt}^{\dagger}\left(\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)\right]\right)^{1/2}\geq$

### v3

@@ 3. $p$-Wasserstein distances and divergences
where $\Pi_{*}^{(\alpha,\beta)}$ stands for a $C_{\mathcal{A},p}$-optimal coupling of the states $\alpha$ and $\beta.$ Although the definition (25) is highly natural in view of the definition of the quadratic Wasserstein divergence [18],
@@
may happen for $p>2$ as shown in the following.
Proposition 3.
For every $p>2$, there exist a finite-dimensional Hilbert space $\mathcal{H}$, a collection $\mathcal{A}\in\mathbf{FCO}(\mathcal{H})$ with $\{\mathcal{A}\}=1$, and pure states $\rho,\omega\in\mathcal{S}(\mathcal{H})$ such that
$d_{\mathcal{A},p}^{p}(\rho,\omega)<0.$
In particular, the quantity
$D_{\mathcal{A},p}^{p}(\rho,\omega)-\frac{1}{2}\Big(D_{\mathcal{A},p}^{p}(\rho,\rho)+D_{\mathcal{A},p}^{p}(\omega,\omega)\Big)$
need not be nonnegative for $p>2$.
Proof.
Let $\mathcal{H}=\mathbf{C}^{3}$ with orthonormal basis $\{e_{1},e_{2},e_{3}\}$, and define
$Q_{i}:=|e_{i}\rangle\langle e_{i}|,\qquad i=1,2,3.$
For some $b_{1},b_{2}>0$ to be fixed later let
$A:=0\cdot Q_{1}+b_{1}Q_{2}+(b_{1}+b_{2})Q_{3}.$
We take $\mathcal{A}=\{A\}$. By (22), the corresponding cost operator is
$C_{\mathcal{A},p}=\sum_{i,j=1}^{3}|a_{i}-a_{j}|^{p}\,Q_{i}\otimes Q_{j}^{T},$
where $a_{1}=0$, $a_{2}=b_{1}$, and $a_{3}=b_{1}+b_{2}$. Expanding this expression, we obtain
$C_{\mathcal{A},p}$ $=b_{1}^{p}(Q_{1}\otimes Q_{2}^{T}+Q_{2}\otimes Q_{1}^{T})+b_{2}^{p}(Q_{2}\otimes Q_{3}^{T}+Q_{3}\otimes Q_{2}^{T})$
$\quad+(b_{1}+b_{2})^{p}(Q_{1}\otimes Q_{3}^{T}+Q_{3}\otimes Q_{1}^{T}).$
Let
$\rho=|u\rangle\langle u|,\qquad\omega=|v\rangle\langle v|$
be pure states. Pure states admit a unique coupling, the product coupling. Hence, by (26),
$d_{\mathcal{A},p}^{p}(\rho,\omega)=\mathrm{tr}_{\mathcal{H}\otimes\mathcal{H}^{*}}\Big[\Big(\omega\otimes\rho^{T}-\frac{1}{2}(\rho\otimes\rho^{T}+\omega\otimes\omega^{T})\Big)C_{\mathcal{A},p}\Big].$
Writing
$r_{i}:=\langle u,Q_{i}u\rangle=|\langle e_{i},u\rangle|^{2},\qquad s_{i}:=\langle v,Q_{i}v\rangle=|\langle e_{i},v\rangle|^{2},$
a direct computation yields
$d_{\mathcal{A},p}^{p}(\rho,\omega)$ $=b_{1}^{p}\big(r_{1}s_{2}+s_{1}r_{2}-r_{1}r_{2}-s_{1}s_{2}\big)$
$\quad+b_{2}^{p}\big(r_{2}s_{3}+s_{2}r_{3}-r_{2}r_{3}-s_{2}s_{3}\big)$
$\quad+(b_{1}+b_{2})^{p}\big(r_{1}s_{3}+s_{1}r_{3}-r_{1}r_{3}-s_{1}s_{3}\big).$
We now choose
$u=e_{2},\qquad v=\frac{e_{1}+e_{3}}{\sqrt{2}}.$
Then
$(r_{1},r_{2},r_{3})=(0,1,0),\qquad(s_{1},s_{2},s_{3})=\Big(\frac{1}{2},0,\frac{1}{2}\Big),$
and therefore
$d_{\mathcal{A},p}^{p}(\rho,\omega)=\frac{1}{2}\,b_{1}^{p}+\frac{1}{2}\,b_{2}^{p}-\frac{1}{4}\,(b_{1}+b_{2})^{p}.$
Finally, choosing $b_{1}=b_{2}=1$, we obtain
$d_{\mathcal{A},p}^{p}(\rho,\omega)=1-2^{p-2}.$
Consequently,
$d_{\mathcal{A},p}^{p}(\rho,\omega)\geq 0\iff p\leq 2,$
and in particular $d_{\mathcal{A},p}^{p}(\rho,\omega)<0$ for every $p>2$. ∎
@@ 4. Relations between the sets of all possible cost operators for different values of $p$
[…]
@@
So let $0<p<\infty$ and $A\in\mathcal{L}(\mathcal{H})^{sa},\,\#\{\mathrm{spec}(A)\}<\infty$ be fixed, and assume that $A$ admits the spectral decomposition $A=\sum_{r=1}^{n}\lambda_{r}Q_{r}.$ Let $J\subseteq[n]$ where the shorthand $n]:=\left\{1,\dots,n\right\}$ is used. Let us define
@@
$\left|\alpha_{J}\right|^{p}\sum_{r\in[n]}\sum_{s\in[n]}\left|\mathbf{1}_{\left\{r\in J\right\}}-\mathbf{1}_{\left\{s\in J\right\}}\right|Q_{r}\otimes Q_{s}^{T}.$ (33)
@@
[…]
@@
If $A\in\mathcal{L}(\mathcal{H})^{sa}$ and $\{\mathrm{spec}(A)\}\leq 2,$ then $A=\lambda Q+\lambda^{\prime}(I-Q)$ for some $Q\in\mathcal{P}(\mathcal{H})$ and $\lambda,\lambda^{\prime}\in\mathbb{R}.$ Therefore,
[…]
@@
$\left|\lambda-\lambda^{\prime}\right|^{p}\left((I-Q)\otimes Q^{T}+Q\otimes(I-Q)^{T}\right).$
@@
$\left|\mu\right|^{p^{\prime}}\left((I-Q)\otimes Q^{T}+Q\otimes(I-Q)^{T}\right).$
@@
[…]
@@
Let $\mathcal{H}$ be a separable Hilbert space, then for any $p^{\prime}\geq p>0$ and any $A=\sum_{i=1}^{n}\lambda_{i}P_{i}\neq c\cdot I$ observable on $\mathcal{H}$ there exists an observable $B=\sum_{i=1}^{n}\mu_{i}P_{i}$ and an observable $A^{\prime}=\sum_{i=1}^{n}\lambda^{\prime}_{i}P_{i}$ on $\mathcal{H}$, such that $\left\{\mathrm{spec}\left(A^{\prime}\right)\right\}<\#\left\{\mathrm{spec}\left(A\right)\right\}$ and
@@
The following proposition complements Conjecture 9 and asserts that the map $p\mapsto\mathbf{C}_{p}(\mathcal{H})$ is not only monotone increasing but strictly monotone increasing on $1,\infty).$
@@
[…]
@@
[…]
@@
Let $Q_{1}$ and $Q_{2}$ be orthogonal projections over $\mathcal{H}$ such that $Q_{1}+Q_{2}\leq I$. Then if $0\leq\lambda\in\mathbb{R}$, then $\lambda(Q_{1}\otimes Q_{2}^{T}+Q_{2}\otimes Q_{1}^{T})\in C_{\infty}(\mathcal{H})$. To see this consider $A(p):=-\frac{\sqrt[p]{\lambda}}{2}Q_{1}+\frac{\sqrt[p]{\lambda}}{2}Q_{2}$. Then
$C_{\{A(p)\},p}=\lambda(Q_{1}\otimes Q_{2}^{T}+Q_{2}\otimes Q_{1}^{T})+$
$\frac{\lambda}{2^{p}}(Q_{1}\otimes(I-Q_{1}-Q_{2})^{T}+(I-Q_{1}-Q_{2})\otimes Q_{1}^{T}+Q_{2}\otimes(I-Q_{1}-Q_{2})^{T}+(I-Q_{1}-Q_{2})\otimes Q_{2}^{T})$ (74)
and
@@ 5. Triangle inequality is inherited by larger parameter values
$0,\infty)\ni t\mapsto t^{\frac{\min\left\{1/p^{\prime},1\right\}}{\min\left\{1/p,1\right\}}}$ (84)
is subadditive. Subadditivity follows from concavity and from the fact that it maps 0 to itself. By the triangle inequality (77), the obvious monotonicity of (84) and the subadditivity of (84) we get
@@
$\left(D_{\mathcal{A},p}^{\max\left\{p,1\right\}}(\tau,\omega)-\frac{1}{2}\left(D_{\mathcal{A},p}^{\max\left\{p,1\right\}}(\tau,\tau)+D_{\mathcal{A},p}^{\max\left\{p,1\right\}}(\omega,\omega)\right)\right)^{\min\left\{\frac{1}{p^{\prime}},1\right\}}$
@@
$\left(D_{\mathcal{A}^{\prime},p^{\prime}}^{\max\left\{p^{\prime},1\right\}}(\tau,\omega)-\frac{1}{2}\left(D_{\mathcal{A}^{\prime},p^{\prime}}^{\max\left\{p^{\prime},1\right\}}(\tau,\tau)+D_{\mathcal{A}^{\prime},p^{\prime}}^{\max\left\{p^{\prime},1\right\}}(\omega,\omega)\right)\right)^{\min\left\{\frac{1}{p^{\prime}},1\right\}}$
$d_{\mathcal{A}^{\prime},p^{\prime}}(\rho,\tau)+d_{\mathcal{A}^{\prime},p^{\prime}}(\tau,\omega).$ (86)
@@ 6. $p=2$ as a possible threshold for the triangle inequality
$\inf_{\Pi\in\mathcal{C}(\rho,\omega)}\left\{\sum_{k=1}^{K}\left(\mathrm{tr}_{\mathcal{H}}\left[\rho A_{k}^{2}+\omega A_{k}^{2}\right]-2\mathrm{tr}_{\mathcal{H}\otimes\mathcal{H}^{*}}\left[\Pi\cdot A_{k}\otimes A_{k}^{T}\right]\right)\right\}.$ (87)
@@
An important feature of the quadratic Wasserstein distances is that the distance of a state $\rho$ from itself (which may be positive) is always realized by the identity channel — see [17, Corollary 1]. Therefore, the quadratic Wasserstein distance of a state $\rho\in\mathcal{S}\left(\mathcal{H}\right)$ from itself admits the explicit form
@@
For every $0<p<2$ there is $\mathcal{H}$ and $\mathcal{A}\in\mathbf{FCO}(\mathcal{H})$ and $\rho,\tau,\omega\in\mathcal{S}\left(\mathcal{H}\right)$ such that
@@
and the positivity condition $\rho\geq 0$ is equivalent to the Euclidean length of $\mathbf{b}_{\rho}$ being at most $1.$ Now observe that Proposition 17 tells us that
@@ 7. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 20
This section is dedicated to the study of the triangle inequality for quadratic quantum Wasserstein divergences. The main result here is a proof of the triangle inequality using the sole assumption that an arbitrary one of the three states involved is pure. This is a generalization of the main result of [6], where we proved the triangle inequality in the case when the intermediate state is pure. It is important to note that ten month after the publication of the first version of this paper, Melchior Wirth proved the triangle inequality for all quadratic quantum Wasserstein divergences in full generality, that is, without any restriction concerning the states involved [54]. His technique relies on advanced complex analysis, and hence is rather different from the argument presented below.
@@
$\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\tau^{1/2}A_{k}\right)^{2}+\left(\omega^{1/2}A_{k}\right)^{2}-2\tau^{1/2}A_{k}\tau^{1/2}\Psi_{opt}^{\dagger}\left(A_{k}\right)\right]\right)^{1/2}\geq$
@@
$\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\tau^{1/2}A_{k}\right)^{2}+\left(\tau^{1/2}\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)^{2}-2\tau^{1/2}A_{k}\tau^{1/2}\Psi_{opt}^{\dagger}\left(A_{k}\right)\right]\right)^{1/2}=$
@@
is a special case of the monotonicity version of Lieb’s concavity — see, e.g., [13, Theorem 1]. We can use (113) for a further estimation in the following way:
@@
$\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\rho^{1/2}\left(\Phi_{opt}^{\dagger}\left(A_{k}\right)-\Phi_{opt}^{\dagger}\left(\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)\right)\right)^{2}\right]\right)^{1/2}\geq$
@@
$\left(\sum_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}\left[\left(\rho^{1/2}A_{k}\right)^{2}+\left(\rho^{1/2}\Phi_{opt}^{\dagger}\left(\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)\right)^{2}-2\rho^{1/2}A_{k}\rho^{1/2}\Phi_{opt}^{\dagger}\left(\Psi_{opt}^{\dagger}\left(A_{k}\right)\right)\right]\right)^{1/2}\geq$